OpenStudy (anonymous):

Verify: 1- sin^2X/1+cosX =CosX

OpenStudy (anonymous):

@abb0t Help!

OpenStudy (abb0t):

\(\frac{ 1-\sin^2(x) }{ 1+\cos(x) }= \cos(x)\) is that correct?

OpenStudy (anonymous):



OpenStudy (abb0t):

Combine denominators. You get: \[\frac{ 1+\cos(x) }{ 1+\cos(x) }-\frac{ 1-\sin^2(x) }{ 1+\cos(x) }\]

OpenStudy (anonymous):

taking lcm we will get (1 + cosx - sin^2x)/1 + cos x now 1 - sin ^2 x is cos^2x just substitute and u will get the answer

OpenStudy (abb0t):

Note: \(1-\sin^2(x)=\cos(x)\)

OpenStudy (abb0t):

Can you figure it out from here?

OpenStudy (anonymous):

No im lost big time>.< wait so who's right?

OpenStudy (anonymous):

Help!D: i dont get this subject at all like at all

OpenStudy (anonymous):

no 1 - sin^2x is cos^2x so the equation becomes (cos^2x + cosx)/1+cosx take cosx common from the numerator u will get cosx(1+cosx)/1+cosx = cosx thus proved :D

OpenStudy (abb0t):

@dhanamkoilraj where did you get the "+' from on ur numerator?

OpenStudy (anonymous):

becoz the numerator is 1 + cosx - sin^2x = cos^2x + cosx

OpenStudy (anonymous):

is # 2

OpenStudy (anonymous):

yea this is the solution to tat question

OpenStudy (anonymous):

This is correct no?

OpenStudy (anonymous):

@abb0t @dhanamkoilraj ?

OpenStudy (anonymous):

So it this correct?

OpenStudy (anonymous):

yea its correct

OpenStudy (anonymous):

Thanks for the approval! ^-^

OpenStudy (anonymous):

no probs :D

OpenStudy (anonymous):

\[CosX \div 1-Sin ^{2}X = SecX\]

OpenStudy (anonymous):

again 1- sin^2x is cos^2x so the equation becomes cosx/cos^2x which is equal to 1/cosx = secx

OpenStudy (anonymous):